OpenStudy (anonymous):

this isnt really physics but.. A truck travels at 40 m/s and slows down to 10 m/s in 5 seconds. What is the acceleration of the truck? Show formula, work, and units for each variable. please help

OpenStudy (anonymous):

Its physics, well you will have to use the kinematic equations. the one your are going to want to use is: v final = v initial + a(t final- t initial) you will plug in your values so it looks like: 10 m/s = 40 m/s + a (0- 5) Then solve for a: .25 = a(-5s) a= -0.05 m/s^2

OpenStudy (anonymous):

I did the math wrong, sorry. Instead of dividing 10 by 40 I should have subtracted so it would look like this: 10 m/s = 40 m/s + a(5-0 s) -30 m/s = a(5 s) -6 m/s^2 = a